One end of a copper poker is placed in a fire with temperature of 600°C, while the other end is kept at room temperature (25°). The poker is 1,5 m long and has a radius of 3 x 10-3 m.


1
Expert's answer
2022-03-28T14:24:05-0400

Answer

In this case,

"\frac{dQ}{dt}=kA(T_2-T_1 )/d\\

A=πr^2"

The amount of heat conducted from one end of the poker to the other in 5.0s is given

"Q=\frac{dQ}{dt }∆t\\=((79.5)*π*(0.005)^2 (502-26))/1.2 (5)\\=12 J."